### v2

# Accurate and Diverse LLM Mathematical Reasoning via Automated PRM-Guided GFlowNets

Abstract
Achieving both accuracy and diverse reasoning remains challenging for Large Language Models (LLMs) in complex domains like mathematics. A key bottleneck is evaluating intermediate reasoning steps to guide generation without costly human annotations. To address this, we first introduce a novel Process Reward Model (PRM) trained automatically using Monte Carlo Tree Search coupled with a similarity-based data augmentation technique, effectively capturing step-level reasoning quality. Leveraging this PRM, we then adapt Generative Flow Networks (GFlowNets) to operate at the reasoning step level. Unlike traditional reinforcement learning focused on maximizing a single reward, GFlowNets naturally sample diverse, high-quality solutions proportional to their rewards, as measured by our PRM. Empirical evaluation shows strong improvements in both accuracy and solution diversity on challenging mathematical benchmarks (e.g., +2.59% absolute accuracy on MATH Level 5 for Llama3.2-3B), with effective generalization to unseen datasets (+9.4% absolute on SAT MATH). Our work demonstrates the potential of PRM-guided, step-level GFlowNets for developing more robust and versatile mathematical reasoning in LLMs.
[…]
360
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable progress in various natural language tasks [4, 8], yet achieving robust and reliable mathematical reasoning remains a significant challenge [18, 11]. While LLMs have shown increasing proficiency in solving mathematical problems [6, 46], current approaches often prioritize accuracy on benchmark datasets [11, 41], potentially overlooking other crucial aspects of intelligent reasoning, such as the ability to explore and generate diverse solution strategies [25]. For LLMs to truly excel in mathematical domains and move beyond pattern recognition towards genuine understanding, they must not only arrive at correct answers but also exhibit the capacity to reason through problems in multiple, varied, and insightful ways [45, 38].
Traditional reinforcement learning methods like Proximal Policy Optimization [PPO; 33] have shown promise in improving LLM mathematical reasoning [44]. However, these methods inherently aim to maximize a single reward signal, often leading to the exploitation of a narrow set of solution strategies [51, 25]. This limitation becomes particularly critical when considering the development of robust and generally applicable problem-solving AI systems, where adaptability to novel situations and the exploration of diverse solution spaces are paramount.
Addressing the need for diverse reasoning requires effective guidance at the level of intermediate steps. However, obtaining reliable reward signals for these steps typically involves expensive human annotation. We overcome this limitation by first developing an automatically trained Process Reward Model [PRM; 38]. Our approach uniquely employs Monte Carlo Tree Search [MCTS; 20] combined with a novel similarity-based data augmentation technique to generate high-quality step-level reward signals without manual labeling. Given this automated and nuanced step-level reward mechanism provided by the PRM, we then propose leveraging Generative Flow Networks [GFlowNets; 1, 2] for fine-tuning. Unlike traditional reinforcement learning methods that often collapse to a single strategy, GFlowNets are designed to sample proportionally to rewards, naturally fostering solution diversity. Our key adaptation is operating GFlowNets at the reasoning step level rather than the token level – each state represents a partial solution, and actions generate complete reasoning steps – allowing the PRM reward to guide the exploration of diverse, high-quality reasoning paths effectively. Prior work using GFlowNets for LLM fine-tuning [13, 36] often utilized variants of the Subtrajectory Balance (SubTB) loss [21], which we adapt for our step-level framework.
Our main contributions are:
  • An efficient and scalable training methodology for mathematical reasoning, featuring an automatically trained Process Reward Model [PRM; 38] guided by Monte Carlo Tree Search [MCTS; 20] and a novel similarity-based data augmentation technique, effectively eliminating the need for expensive human annotations.
  • A step-level GFlowNet framework for mathematical reasoning, where states and actions correspond to meaningful solution components rather than individual tokens, enabling fine-grained control over reasoning path generation.
  • Comprehensive empirical validation showing both improved accuracy and enhanced solution diversity. Our approach achieves a 17.05% accuracy on MATH Level 5 using Llama3.2-3B (compared to 14.46% baseline and 15.32% with PPO), while generating measurably more diverse solution strategies. Additionally, our method generalizes to datasets unseen during the fine-tuning phase. For instance, we observe a relative increase of $14.32\%$ is observed on the SAT MATH dataset. Similar improvements are observed with an 8B model.
Empirically, we demonstrate that our approach not only improves accuracy on challenging mathematical reasoning benchmarks but also generates more diverse solution strategies compared to both baseline models and PPO-fine-tuned variants. This is particularly evident in our diversity analysis, where GFlowNet-fine-tuned models show significantly lower semantic similarity between generated solutions while maintaining correctness.
Our work points towards new directions for developing next-generation LLMs with enhanced reasoning capabilities. By demonstrating a method to improve both accuracy and diversity in a complex domain like mathematical reasoning, we open possibilities for creating more robust, versatile, and ultimately more intelligent LLM systems capable of tackling a wider range of challenging problems.
2 Background and Related Work
Mathematical Reasoning in LLMs has seen significant progress through various approaches, including chain-of-thought prompting [42], self-consistency [41], and reward modeling [19, 48]. While these methods have improved accuracy, they often lack mechanisms for promoting solution diversity. Recent work by Wang et al. [40] highlights the importance of diverse reasoning paths but focuses primarily on accuracy rather than explicitly encouraging diversity during training.
Chain-of-thought (CoT) approaches [42] and their variants such as Tree-of-Thought [43] and Graph-of-Thought [3] have demonstrated substantial improvements by encouraging models to articulate intermediate steps in their problem-solving process. However, the precise mechanisms underlying these improvements remain an active area of investigation, with some researchers questioning whether the benefits derive specifically from human-like task decomposition or simply from the additional computation afforded by generating more tokens [29]. Several enhancements to CoT approaches have been proposed, including using datasets of preference pairs of reasoning traces to finetune the CoT-generating model [17].
Despite these advances, current approaches to mathematical reasoning predominantly emphasize accuracy improvements rather than fostering diverse solution strategies. This limitation becomes particularly relevant in educational contexts and mathematical exploration, where multiple valid approaches can provide deeper insights into problem structures. This gap motivates our investigation into leveraging generative flow networks to promote both accuracy and diversity in mathematical reasoning.
Generative Flow Networks (GFlowNets) represent a novel framework for learning to sample from a desired distribution, offering advantages over traditional reinforcement learning approaches [1, 2]. GFlowNets operate on a directed acyclic graph (DAG) structure, where states $\mathcal{S}$ represent partial constructions and actions $\mathcal{A}$ represent transitions between states. This graph contains a unique source state $s_{0}$ with no parents and a sink state $s_{f}$ with no children. States that connect directly to $s_{f}$ are termed terminal states $\mathcal{X}$, each associated with a positive reward $R(x)>0$ for $x\in\mathcal{X}$.
The core objective of GFlowNets is to learn policies that generate complete trajectories $\tau=(s_{0},s_{1},...,s_{n},s_{f})$ such that terminal states are sampled with probabilities proportional to their rewards: $P(x)\propto R(x)$. This is achieved by learning a flow function $F:\mathcal{A}\rightarrow\mathbb{R}^{+}$ satisfying two fundamental constraints:
  1. 1.
    Flow conservation: For every non-terminal, non-source state $s\in\mathcal{S}\setminus\{s_{0},s_{f}\}$:
    $\sum_{(s^{\prime},s)\in\mathcal{A}}F(s^{\prime}\rightarrow s)=\sum_{(s,s^{\prime\prime})\in\mathcal{A}}F(s\rightarrow s^{\prime\prime})$ (1)
  2. 2.
    Reward matching: For every terminal state $x\in\mathcal{X}$:
    $F(x\rightarrow s_{f})=R(x)$ (2)
From this flow function, forward and backward policies can be derived as:
$P_{F}(s^{\prime}|s)=\frac{F(s\rightarrow s^{\prime})}{F(s)}\text{ and }P_{B}(s|s^{\prime})=\frac{F(s\rightarrow s^{\prime})}{F(s^{\prime})}$ (3)
where $F(s)=\sum_{(s,s^{\prime})\in\mathcal{A}}F(s\rightarrow s^{\prime})$ represents the total outflow from state $s$.
Unlike reinforcement learning methods that focus on maximizing expected cumulative rewards - typically converging to deterministic policies that exploit highest-reward paths - GFlowNets learn stochastic policies that maintain diversity while still favoring high-reward solutions. This balance between exploration and exploitation makes GFlowNets particularly valuable for applications where multiple viable solutions are preferable to a single optimal one.
Various training objectives have been proposed to learn effective GFlowNets. The flow-matching objective [1] directly enforces the flow constraints. The detailed-balance (DB) objective [2] operates at the level of individual transitions, enforcing a local equilibrium:
$F(s)P_{F}(s^{\prime}|s)=F(s^{\prime})P_{B}(s|s^{\prime})$ (4)
The trajectory-balance (TB) objective [22] provides a more global perspective by considering complete trajectories:
$\mathcal{L}_{TB}(\tau;\theta)=\left(\log\left(\frac{Z_{\theta}\prod_{t=1}^{n+1}P^{\theta}_{F}(s_{t}\mid s_{t-1})}{R(s_{n})\prod_{t=1}^{n}P^{\theta}_{B}(s_{t-1}\mid s_{t})}\right)\right)^{2}$ (5)
where $Z_{\theta}$ is a learnable parameter representing the total reward. Additional approaches include SubTB($\lambda$) [21], which enables learning from incomplete trajectories, and weighted DB objectives [34] that account for terminal state descendants to improve training dynamics. Such objectives share similarities with hierarchical variational inference algorithms, as described in Malkin et al. [23].
GFlowNets have demonstrated remarkable success across diverse domains. Their ability to generate diverse, high-quality samples addresses a fundamental challenge in scientific discovery tasks involving astronomically large search spaces. By learning to sample diverse high-reward candidates, GFlowNets can efficiently identify promising regions of the design space while maintaining the variety needed to accommodate additional constraints not captured in the primary reward function, such as synthesis feasibility or absence of side effects. They have been successfully applied to molecule generation [14], Bayesian structure learning [7], causal discovery [24, 7], material design [5, 26], drug discovery [28], biological sequence design and editing [14, 9], scheduling [49], and combinatorial optimization problems [47]. Recent theoretical extensions have expanded GFlowNets to continuous spaces [16], stochastic environments [27], and adversarial settings [15].
Recent work has explored GFlowNets for language generation tasks [13, 36, 12], typically operating at the token level. However, token-level approaches can struggle with maintaining semantic coherence in complex reasoning tasks, where the relationship between local token choices and global solution quality can be difficult to capture. Our work addresses this limitation by adapting GFlowNets to operate at the reasoning step level, enabling more controlled and meaningful solution generation while preserving the diversity benefits that make GFlowNets attractive for creative and exploratory tasks.
3 Process Reward Model for Mathematical Reasoning Steps
We train a PRM $U(s^{\prime}|s)$ that evaluates the quality of a proposed reasoning step $s^{\prime}$ given its predecessor $s$. The PRM outputs a score between 0 and 1, representing the probability that step $s^{\prime}$ will lead to a correct solution, assuming $s$ is correct. Unlike previous work that uses binary labels [20, 19, 40], our continuous scores enable more nuanced evaluation of step quality.
3.1 Automated Data Generation via MCTS
To train the PRM without human annotations, we use MCTS coupled with a novel data augmentation strategy, based on rollout reuse and similarity grouping (details provided in subsections 3.1, 3.2, and 3.3). For each candidate step, we perform $k=96$ rollouts (similar to Luo et al. [20]) to estimate its Monte Carlo value $MC(s)$. Crucially, we only include steps in our training dataset up to the first incorrect step (where $MC(s)=0$), as subsequent steps would be built upon invalid reasoning. The steps are identified using line breaks, and the model used to generate the data is prompted to separate the steps in its reasoning using 4 few-shots prompts.
We use a binary search strategy to efficiently identify the first incorrect step within a candidate reasoning path; subsequent steps originating from this flawed logic are disregarded during evaluation. To assess the quality of a specific step $s$ identified during this search, we perform $k$ Monte Carlo rollouts ($k=96$ in our experiments to ensure representative sampling). For each rollout, the LLM generates a complete solution attempt with a temperature of 0.6 for diversity, prompted with the original question and the solution prefix ending just before step $s$. The value of the step, denoted $MC(s)$, is then calculated as the proportion of these $k$ rollouts that ultimately yield the correct final answer to the problem.
$MC(\text{s})=\frac{\text{number of successful rollouts}}{\text{total number ofrollouts}}$
All the rollouts are then placed in a pool with statistics N(s), MC(s),Q(s,r). Which are defined as follow :
  1. 1.
    N(s) : Number of visit of state s
  2. 2.
    Q(s,r) : state-rollout value function defined as
    $Q(\text{s,r})=\alpha^{1-MC(\text{s})}\beta^{\frac{\text{len(r)}}{L}}$
    where $\alpha,\beta\in(0,1]$ and $L>0$ are hyperparameters
This process continues until the first error in the reasoning has been detected. Once this has been done, we can then select a new rollout among all the rollouts available in the pool using their statistics and repeat the process, creating a tree. Rollouts $(s,r)$ are selected according to tree statistics : $(s,r)=argmax_{(s,r)}[Q(s,r)+U(s)]$ as defined by [20], which is based on a variant of the PUCT algorithm [32], and where:
$U(s)=c_{puct}\frac{\sqrt{\sum_{i}{N(s_{i})}}}{1+N(s)},$
where $c_{puct}$ is a constant determining the level of exploration. This strategy initially favors rollouts with low visit counts but gradually shifts preference towards those with high rollout values. Then, as we go along, we will sample rollouts that have larger and larger values in accordance with Q(s,r), This is advantageous because instead of sampling the rollouts uniformly, this strategy allows us to sample rollouts that have larger MC(s) values, which means that the stage is of fairly good quality. This strategy would make the PRM more efficient because a PRM capable of detecting errors in good quality reasoning will be more useful than a PRM capable of detecting errors in reasoning that is initially poor.
The algorithm stops if there are no more rollouts available in the pool or if a predefined limit is reached.
0:  Question $q$, Language Model LM, Number of completions $k=96$, Temperature $T=0.6$
0:  Dataset of training examples
1:  Initialize root state $r_{\text{root}}\leftarrow q$
2:  Initialize tree with root node $s_{\text{root}}$ containing $r_{\text{root}}$
3:  Initialize visit count $N(s_{\text{root}})\leftarrow 0$
4:  Initialize Monte Carlo estimation $MC(s_{\text{root}})\leftarrow 0$
5:  Initialize state-rollout value function $Q(s_{\text{root}},r)\leftarrow 0$
6:  while not converged do
7:     Select a trajectory using PUCT algorithm based on $Q(s,r)$ and $U(s)$
8:     Perform binary search to locate the first incorrect step in the selected trajectory, the step currently sampled is $s_{\text{candidate}}$
9:     Generate $k$ completions using temperature sampling from $\text{LM}(c|s_{\text{candidate}})$
10:     Initialize $correct\_count\leftarrow 0$
11:     for each completion $c_{i}$ do
12:        if rollout is correct then
13:           $correct\_count\leftarrow correct\_count+1$
14:        end if
15:     end for
16:     $MC(s_{\text{candidate}})\leftarrow\frac{correct\_count}{k}$
17:     if $correct\_count=0$ then
18:        Add $s_{\text{candidate}}$ to the dataset with value 00
19:        This step is considered to be incorrect
20:        Break
21:     else
22:        Update tree by adding $s_{\text{candidate}}$ with value $MC(s_{\text{candidate}})$ and add all the generated rollouts to the tree as well as the number of visits of the node $N(s_{\text{candidate}})$ and $Q(s_{\text{candidate}},r)$
23:     end if
24:  end while
25:  Dataset Collection Phase:
26:  for each node $s$ in the tree do
27:     Add $s$ to the dataset with its evaluated step, value and all the rollouts generated from this step.
28:  end for
28:  Dataset of training examples
Algorithm 1 MCTS-based data generation
We applied this MCTS data generation process (Algorithm 1) using the Qwen2.5-Math model to an initial set of 70,000 mathematical problems sourced from the OpenMathInstruct2 dataset [37].
3.2 Dataset Augmentation via Rollout Reuse and Similarity Grouping
To improve the efficiency of our data generation and create a larger, more comprehensive dataset, we implement a dataset augmentation strategy. This strategy leverages the rollouts generated during the MCTS process and incorporates a step similarity grouping technique.
During MCTS, for each evaluated step, we store the $k$ rollouts used to estimate its $MC(s)$ value. Each rollout is stored as a tuple $(r,x)$, where $r$ is the generated reasoning path (rollout) and $x$ is a binary indicator (1 or 0) denoting whether the rollout led to a correct final answer. To augment our dataset, we extract individual steps from these stored rollouts. For each rollout step, we create a new data entry. The “prefix” for this new entry is constructed by concatenating the original prefix of the step from which the rollouts were generated with the step itself. This process allows us to reuse steps from successful and unsuccessful rollouts, significantly increasing the size of our dataset. Ideally, for each step evaluated in MCTS, we could add up to $k$ new step examples to our dataset through this rollout reuse.
However, directly adding all rollout steps without further processing can introduce inconsistencies. As illustrated in Table 1, semantically similar steps can receive different $MC(s)$ values if evaluated in isolation with a limited number of rollouts ($k=1$, which would be the case if we directly evaluated each step in a rollout independently). This is because evaluating with $k=1$ can lead to noisy and unreliable step value estimations.
Table 1: Example of Similar Steps with Different Values
 Steps value
Max attended 40 college courses in 2 years 0
Max attended 40 college courses 1
To address this consistency issue and further refine our dataset, we introduce a post-processing step based on step similarity grouping. First, we define a Step Similarity Function to quantify the similarity between two steps of mathematical reasoning. This function evaluates similarity based on two primary criteria: (1) Calculation Consistency: If both steps contain mathematical calculations, the function checks if the results of these calculations are identical. If the results differ, the steps are considered dissimilar, and the function returns a similarity score of 0. (2) Textual Similarity: If the calculation results are the same, or if neither step contains calculations, the function computes the Levenshtein distance between the textual content of the two steps.
If the similarity score computed by the Step Similarity Function exceeds a predefined threshold (set to 0.85 in our experiments), the steps are grouped into the same similarity group. Within each group, we then assign consistent step values. Specifically:
  • If all steps within a group originated from rollouts that led to correct final answers, all steps in the group are assigned a value of 1.
  • If all steps within a group originated from rollouts that led to incorrect final answers, all steps in the group are assigned a value of 0.
  • If there is a mixture of correct and incorrect rollouts associated with the steps in a group, all steps in the group are assigned the $MC(s)$ value that was originally estimated for the MCTS step from which these rollouts were generated. In this case, we use the original $MC(s)$ as a reasonable approximation of the value for all similar steps in the group, assuming that steps similar to a step with a known $MC(s)$ are likely to have a similar probability of leading to a correct solution.
This combined dataset augmentation process ensures greater consistency and reliability in the step-level labels. The initial MCTS generation yielded approximately 100k step examples from the input problems described in Section 3.1. Applying rollout reuse and similarity grouping significantly expanded this set to the final 2.1M entries used for PRM training, greatly enhancing dataset size and diversity and improving the PRM’s generalization capability. This final dataset comprises approximately 30% false steps, 20% steps guaranteed correct (value 1), and 50% intermediate steps (value between 0 and 1). The complete data processing workflow, including the handling of incorrect steps during MCTS and the augmentation logic, is visually summarized in Figure 1.
Figure 1: Data processing workflow for PRM training. Starting from a step $s$ with $MC(s)=1/2$, the diagram shows how subsequent steps are processed based on their Monte Carlo values. Similar steps (indicated by dashed boxes) share MC values. Steps following an incorrect step ($MC=0$) are excluded from the training dataset, as they would be built upon invalid reasoning. Gray boxes indicate steps that become irrelevant to the training process.
3.3 PRM Training
We fine-tuned Qwen2.5-Math [30] as our PRM. We train the PRM to predict the probability of a step leading to a correct solution using Binary Cross-Entropy Loss (BCELoss):
$L=-\frac{1}{N}\sum_{i=1}^{N}\left[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right],$
where $y_{i}$ is the true label for the $i$-th step (i.e., the $MC(s)$ value), and $\hat{y}_{i}$ is the predicted probability for the $i$-th step, given by:
[…]
Training was conducted for 1 epoch with a learning rate of 8e-6, batch size 144, and weight decay 0.01 on 8 A100 GPUs.
3.4 PRM Validation
We validate our PRM through three complementary analyses. First, we assess its ability to detect step-level errors by corrupting correct solution steps in systematic ways (e.g., changing numbers, or removing key reasoning components). The PRM consistently assigns lower scores to corrupted steps, demonstrating its sensitivity to reasoning errors.
Second, we evaluate its capability to support diverse solution paths by comparing scores assigned to different valid approaches for the same problem. The PRM assigns comparable scores (within 0.1) to different valid approaches, indicating its ability to recognize multiple correct reasoning paths. Detailed validation results and example cases are provided in subsections 3.4.1 and 3.4.2.
Thirdly, we implemented a guided search strategy to enhance the mathematical reasoning capabilities of LLMs using our trained PRM, similar to the approach in Snell et al. [35]. At each step of generation, the LLM proposes $k$ candidate steps (generated with a temperature of 0.8 for diversity), and the PRM selects the step with the highest score. This step is then appended to the prompt, and the process is repeated until a complete solution is generated.
Figure 2: Accuracy of Falcon models on a subset of MATH Hard using PRM-guided search with varying numbers of proposed steps ($k$). Dashed horizontal lines indicate the baseline accuracy of unguided Falcon3-3B, 7B, and 10B models using prompt-based decoding. Curves represent the accuracy of Falcon3-1B, 3B, and 7B using PRM-guided search with different $k$ values. This demonstrates the effectiveness of PRM-guided search in improving LLM accuracy on challenging mathematical reasoning tasks, especially for smaller models.
Figure 2 demonstrates the effectiveness of this PRM-guided search on a subset of the MATH Hard dataset. The figure compares the accuracy of different Falcon models (1B, 3B, and 7B) using PRM-guided search with varying numbers of proposed steps ($k$) against the baseline accuracy of larger, unguided models using prompt-based decoding.
These results highlight the effectiveness of our PRM in guiding the step-by-step reasoning of LLMs. By selecting the most promising steps based on the PRM’s evaluation, the guided search strategy significantly enhances the accuracy of smaller models, even allowing them to surpass the performance of larger, unguided models on challenging mathematical reasoning tasks. Similar use of PRMs can be found in Wang et al. [39, 40] or other “helper” models such as Preference Process Model in Guan et al. [10]
3.4.1 Error Detection Validation
We systematically corrupted correct solution steps in three ways: Number manipulation, Logic corruption, and Step reordering. The PRM showed consistent score reductions.
3.4.2 Solution Diversity Validation
We evaluated PRM scores on MATH problems with multiple solutions. The PRM assigned comparable scores to different valid reasoning steps, indicating support for multimodal reasoning as can be seen in Figure3.
Figure 3: Reasoning steps and corresponding PRM scores. Valid steps from different approaches receive high, comparable scores, while corrupted steps receive lower scores.
4 Step-Level GFlowNet Fine-tuning for Diverse Solutions
4.1 Framework definition
Building on Hu et al. [13], we adapt GFlowNets to operate at the reasoning step level rather than the token level. In our framework, a state $s$ represents a partial solution (the question and generated reasoning steps), and an action corresponds to appending an extra step to $s$ to obtain another partial solution $s^{\prime}$. This step-level granularity allows more semantically meaningful control over solution generation.
The forward policy $\pi(s^{\prime}|s)$, parameterized by an LLM, defines the probability of generating step $s^{\prime}$ given state $s$. Following Bengio et al. [2], we incorporate a sink state $s_{f}$ to handle variable-length solutions, with $\pi(s_{f}|s)$ representing the termination probability. The reward $R(s_{n})$ is only defined for states that can transition to $s_{f}$ (terminating states) and is computed as:
$R(s_{n})=\prod_{i=1}^{n}U(s_{i}|s_{i-1}).$ (6)
During generation, we use temperature sampling ($T=0.6$) to promote diversity in the proposed steps. This is particularly important as GFlowNets are designed to sample from the reward-proportional distribution rather than simply maximizing rewards.
4.2 Training Objective and Implementation
We use the modified SubTB loss [21] of Hu et al. [13], adapted for our step-level setting:
$\mathcal{L}=\sum_{0\leq i<j\leq n}\lambda^{j-i}\left(\log\frac{R(s_{i})\prod_{k=i+1}^{j}\pi(s_{k}|s_{k-1})\pi(s_{f}|s_{j})}{R(s_{j})\pi(s_{f}|s_{i})}\right)^{2}$ (7)
Our training procedure maintains a replay buffer $\mathcal{B}$ of size 1000 storing complete trajectories. The replay buffer prioritizes trajectories with higher rewards to accelerate learning. For each question, we:
  1. 1.
    Generate $k$ candidate reasoning paths using current policy
  2. 2.
    Evaluate rewards for terminating states using PRM
  3. 3.
    Update policy using modified SubTB loss
  4. 4.
    Store successful trajectories in $\mathcal{B}$
0:  Question $q$, Policy Model $\pi_{\theta}$ (LLM), PRM $U$, Replay Buffer $\mathcal{B}$, Generations per question $k$, Temperature $T$, Batch Size $B$
1:  for each batch of questions in the training dataset do
2:     for each question $q$ in the batch do
3:        Generate $k$ responses for question $q$ using $\pi_{\theta}$ with temperature $T$
4:        Split each response into steps and evaluate reward $R(s_{1:i}s_{f})$ for each trajectory using the PRM, storing trajectories in $\mathcal{B}$
5:        Sample a batch of size $B$ from the replay buffer $\mathcal{B}$
6:        for  each trajectory in the batch do
7:           Compute the loss $\mathcal{L}$ based on the reward function $R$ and policy model $\pi_{\theta}$
8:        end for
9:        Perform one step of optimization to minimize the loss $\mathcal{L}$ with respect to the parameters $\theta$ of $\pi_{\theta}$
10:     end for
11:  end for
11:  Trained policy model $\pi_{\theta}$ with GFlowNets
Algorithm 2 GFlowNet-finetuning
To further improve training efficiency and focus learning on promising reasoning paths, we implemented a prioritized replay mechanism in our replay buffer. Instead of uniformly sampling trajectories from the replay buffer, we prioritized trajectories that had achieved higher rewards, as evaluated by the PRM. This prioritization ensures that the GFlowNet policy is preferentially trained on experiences that have been deemed more successful and of higher quality, accelerating learning and improving convergence.
5 Experiments
For the main fine-tuning experiments, we used Llama3 models as the base architectures (specific sizes detailed in subsection 5.1). These models were fine-tuned using either our step-level GFlowNet approach or a PPO baseline. Both fine-tuning methods utilized reward signals from our Process Reward Model (PRM), which was pre-trained using the methodology described in Section 3 (based on the Qwen2.5-7B-math model [30]). The GFlowNet and PPO fine-tuning was conducted on a challenging subset of 10,000 questions from the OpenMathInstruct2 dataset [37] (“augmented math” category, aligning with MATH Level 5 difficulty). We evaluated the performance of the fine-tuned models on the MATH Hard benchmark (Level 5 subset) [11], the GSM8K benchmark [6], and the SAT MATH dataset [50] to assess both in-domain accuracy and generalization capabilities.
5.1 Experimental Setup
We conduct experiments on Llama3 [8] using two model sizes:
  • Llama3.2-3B-it: A smaller model to demonstrate efficiency
  • Llama3.1-8B-it: A medium-sized model for performance comparison
Hyperparameters:
  • Learning Rate: We used a learning rate of 5e-6 for the Adam optimizer during GFlowNet fine-tuning. To optimize the learning rate schedule, we used a cosine scheduler, which gradually decreases the learning rate over the course of the training epoch.
  • Batch Size: A batch size of 144 trajectories was used for each training iteration. This batch size was chosen to balance computational efficiency and the stability of gradient updates.
  • Gradient Clipping: To prevent exploding gradients during training, we applied gradient clipping with a maximum norm of 1.0. This technique helps to stabilize the training process, particularly in the context of recurrent neural networks like the LLMs used in our GFlowNet policy.
  • $\lambda$ Value for SubTB Loss: The $\lambda$ hyperparameter in the Subtrajectory Balance (SubTB) loss function controls the discount factor for subtrajectory rewards. We set $\lambda=1.0$ in our experiments. This value implies no discounting, giving equal weight to all subtrajectory balance terms in the loss.
  • Replay Buffer Size: To stabilize training and improve sample efficiency, we utilized a replay buffer of size 1000. This buffer stores previously generated complete trajectories, allowing the GFlowNet policy to learn from a diverse set of high-reward experiences.
5.2 Main Results
Table 2 presents our main results, comparing GFlowNet fine-tuning against PPO and baseline models:
Table 2: Performance Comparison on Mathematical Reasoning Tasks
Model MATH Level 5 sat math GSM8K
Llama3.2-3B-it 14.46% 65.6% 67.8%
+ PPO 15.32% 70.0% 68.4%
+ GFlowNet 17.05% 75.0% 68.5%
Llama3.1-8B-it 17.96% 81.2% 78.1%
+ PPO 18.44% 81.2% 79.1%
+ GFlowNet 18.67% 84.4% 79.0%
The empirical results validate the effectiveness of step-level GFlowNet fine-tuning across model scales. The smaller Llama3.2-3B-it model shows substantial improvements with GFlowNet fine-tuning, achieving a 2.59 percentage point gain over the baseline on MATH Level 5. Similarly, the larger Llama3.1-8B-it model demonstrates consistent improvements, with particularly striking gains on the SAT MATH dataset where GFlowNet fine-tuning achieves an 85% accuracy - a significant 3.8 percentage point improvement over both baseline and PPO variants. These findings demonstrate that GFlowNets can enhance mathematical reasoning capabilities regardless of model size, while showing remarkable generalization to tasks beyond the training domain.
Importantly, while PRM training has a computational cost, our analysis, detailed in subsection 5.4, shows that GFlowNet fine-tuning itself is surprisingly efficient, requiring less training time than comparable PPO baselines.
5.3 Solution Diversity Analysis
To quantitatively evaluate the diversity of solutions generated by GFlowNet fine-tuning, we used a semantic similarity metric. We measured the pairwise semantic similarity between the reasoning steps generated by each model using the “paraphrase-MiniLM-L6-v2” model [31], a pre-trained sentence embedding model. Each step was encoded into a high-dimensional space using the encoding function of the model, with the output stored as tensors. The pairwise similarity between these embeddings was then computed using the cosine similarity measure, a standard metric for assessing vector similarity in semantic spaces. This metric provides a continuous score representing the semantic proximity between different reasoning steps, where lower scores indicate greater dissimilarity and, consequently, higher solution diversity. The average semantic similarity across multiple generated solutions for each model is presented in Table 3.
Table 3: Solution Diversity Analysis
Model Semantic Similarity
Llama3.2-3B-it 0.80
+ PPO 0.82
+ GFlowNet 0.78
The semantic similarity scores presented in Table 3 provide compelling evidence that GFlowNet fine-tuning effectively enhances the diversity of generated mathematical solutions. This consistent reduction in semantic similarity indicates that GFlowNet-fine-tuned models are indeed generating reasoning paths that are more semantically distinct and varied compared to the other approaches. This enhanced diversity is a direct consequence of the GFlowNet training objective, which, unlike reward-maximizing RL methods like PPO, is explicitly designed to sample from a distribution proportional to the reward.
5.4 Training Efficiency
To assess the computational efficiency of our GFlowNet fine-tuning approach, we conducted a comparative analysis of the training time and computational resources required for PRM training, GFlowNet fine-tuning, and PPO baseline training. These experiments were performed using consistent hardware and training data settings to ensure a fair comparison. The results of this training efficiency comparison are summarized below:
Computational Resources: All training experiments were conducted on a cluster of machines equipped with NVIDIA A100 GPUs. For PRM training, we utilized 8 A100 GPUs in parallel to accelerate the data generation and model fine-tuning process. GFlowNet and PPO fine-tuning experiments were conducted using the same hardware setup for consistent resource allocation.
Training Time Comparison:
  • PRM Training: Training the Process Reward Model (PRM), including the automated data generation phase using MCTS and the subsequent PRM fine-tuning, required approximately 4 hours of training time using 8 A100 GPUs. The data generation phase using MCTS constitutes a significant portion of this training time.
  • GFlowNet Fine-tuning: Fine-tuning a GFlowNet policy for a specific LLM (e.g., Llama3.2-3B-it or Llama3.1-8B-it) using our step-level approach typically required around 1 hour of training time on the allocated hardware for 10,000 questions from the OpenMathInstruct2 dataset [37].
  • PPO Baseline Training: Training the PPO baseline models, using the same PRM for reward guidance and with comparable hyperparameter settings, generally required approximately 2 hours of training time for 10,000 questions from the OpenMathInstruct2 dataset [37]. This is longer than the GFlowNet fine-tuning time, potentially indicating a greater sample efficiency or faster convergence of the GFlowNet training process in our setup.
6 Conclusion
We have introduced a novel step-level GFlowNet framework for mathematical reasoning that achieves two crucial objectives: improving accuracy and promoting solution diversity. Our approach demonstrates that operating at the reasoning step level, rather than the token level, enables more effective control over the generation process while maintaining semantic coherence.
The empirical results show significant improvements over both baseline models and PPO fine-tuning, particularly for smaller models. This suggests that our approach could be especially valuable in resource-constrained settings where smaller models are preferred. Furthermore, the increased solution diversity achieved by our method aligns well with educational applications, where exposure to multiple valid solution strategies can enhance learning outcomes.
Limitations
Despite these promising results, several limitations warrant discussion, it is noteworthy that the computational cost of initial PRM training through MCTS remains significant, and that the approach requires careful tuning of the replay buffer strategy
Future Work
Promising directions for future research include exploring offline GFlowNet training to reduce computational costs, developing more sophisticated semantic diversity metrics, and investigating the educational impact of diverse solution generation